# Individualizing Glioma Radiotherapy Planning by Optimization of a Data and Physics-Informed Discrete Loss

### v3

@@ 4.7 Initial Guess
We aim to solve the optimization problem for the model parameters referenced in a table in Figure 1 as well the tumor concentration field $u(x,y,z,t)$ on a 4D discrete grid. A meaningful initial guess for these values is crucial for the time of the optimization process and the overall success. We assume the initial tumor location coordinates $x_{0},y_{0},z_{0}$ to be in the center of the tumor core. In addition, for the initial guess we assume $\{R,\theta_{\text{PET}},\theta_{\text{BKG}},\theta_{\text{down}},\theta_{\text{up}}\}$ to be in the middle of the plausible range[30]. Here, we describe the procedure followed to obtain the remaining $\{u,D,\rho\}$:
@@
4.8 Time Complexity
The execution times of the studied methods are summarized as follows. The model $\text{PDE}_{\text{CMA-ES}}$ exhibited an evaluation time of $80\pm 55$ minutes per patient. For GliODIL and its associated $\text{PDE}_{\text{GliODIL}}$, the patient evaluation time was significantly lower, at $45\pm 20$ minutes. The $\text{PDE}_{\text{LMI}}$ model demonstrated the quickest evaluation time, taking only 5 minutes. It’s important to note that the evaluations for GliODIL and $\text{PDE}_{\text{GliODIL}}$ were conducted on a single GPU, whereas the computations for $\text{PDE}_{\text{CMA-ES}}$ and $\text{PDE}_{\text{LMI}}$ were performed on 16 CPU cores.
5 Data and Code Availability
The synthetic data generated for this study, along with the 152 real patient data utilized (FET-PET and segmentations), are both available for reproduction and benchmarking using the resources provided at github.com/m1balcerak/GliODIL/.
@@ 6 Acknowledgment
We would like to extend our gratitude to Bastian Wittmann from the University of Zurich and Tobias Duswald from the Technical Universtity of Munich for their assistance with the manuscript.

### v4

@@ 4.8 Initial Guess
We aim to solve the optimization problem for the model parameters referenced in a table in Figure 1 as well the tumor concentration field $u(x,y,z,t)$ on a 4D discrete grid. A meaningful initial guess for these values is crucial for the time of the optimization process and the overall success. We assume the initial tumor location coordinates $x_{0},y_{0},z_{0}$ to be in the center of the tumor core. In addition, for the initial guess we assume $\{R,\theta_{\text{PET}},\theta_{\text{BKG}},\theta_{\text{down}},\theta_{\text{up}}\}$ to be in the middle of the plausible range[Harpold2007]. Here, we describe the procedure followed to obtain the remaining $\{u,D,\rho\}$:
@@
We also investigate how the initial guess affects solution quality by varying $D_{\text{ch}}$, $\rho_{\text{ch}}$, $x_{0}$, $y_{0}$, and $z_{0}$ by ±20% using a uniform distribution. For real patient data (patient AA), the Core Dice coefficient was $0.81\pm 0.02$, edema Dice $0.35\pm 0.05$, and PET core Dice $0.65\pm 0.05$, all achieved without changes in convergence time. These results indicate that reasonable variations in the initial guess only modestly affect segmentation fit quality. However, the initial guess remains important, as lacking it or using overly simplistic guesses (e.g., duplicating the initial condition across temporal resolution) can prevent the model from converging to meaningful solutions. Nonetheless, as long as the initial guess is reasonable, variations make little difference.
4.9 Execution Times and Spatial Resolutions
The execution times of the studied methods are summarized as follows. The model $\text{PDE}_{\text{CMA\_ES}}$ exhibited an evaluation time of $80\pm 55$ minutes per patient. For GliODIL and its associated $\text{PDE}_{\text{GliODIL}}$, the patient evaluation time was significantly lower, at $45\pm 20$ minutes, while nnU-Net was even faster, taking only a few seconds. It’s important to note that the evaluations for GliODIL, $\text{PDE}_{\text{GliODIL}}$, and nnU-Net were conducted on a single GPU, whereas the computations for $\text{PDE}_{\text{CMA\_ES}}$ and $\text{PDE}_{\text{LMI}}$ were performed on 16 CPU cores.
In terms of spatial resolution, $\text{PDE}_{\text{CMA\_ES}}$ employs a forward Euler finite difference scheme with a progressively increasing spatial resolution, reaching a maximum of $96^{3}$ for the entire space. $\text{PDE}_{\text{GliODIL}}$, on the other hand, uses a Crank-Nicolson scheme with a resolution of $48^{3}$ for the tumor region, resulting in a comparable total spatial coverage. However, since the PDE in use is nonlinear, there are no guarantees regarding scaling the temporal resolution needed for stability at higher spatial resolutions. In contrast, $\text{PDE}_{\text{LMI}}$ operates on a fixed $128^{3}$ resolution, consistent with its training scale.
5 Data Availability
Source data are provided with this paper. The patient data utilized (including tissues, FET-PET scans, and segmentations) are available for reproduction and benchmarking at https://huggingface.co/datasets/m1balcerak/GliODIL.
6 Code Availability
The code for the proposed method, as well as for generating synthetic patients, can be found at https://github.com/m1balcerak/GliODIL.
@@ 7 Acknowledgment
We extend our gratitude to Bastian Wittmann (University of Zurich) for his assistance with the manuscript. We also thank the Dagstuhl Seminar (2023) for fostering discussions that inspired this work. This research was supported by the European High-Performance Computing Joint Undertaking (EuroHPC) grant DCoMEX (956201-H2020-JTI-EuroHPC-2019-1), the Helmut Horten Foundation, the European Cooperation in Science and Technology (COST), and the National Science Foundation (NSF), Division of Mathematical Sciences (DMS), under grants NSF-DMS-1763272/Simons Foundation (QN594598, JL) and NSF-DMS-2309800 (JL).
8 Author Contributions
M.B. was responsible for the methodology, software implementation, and writing the original draft. J.W. contributed to the methodology and the implementation of baselines. P.K. handled the methodology and software implementation. I.E. was involved in the methodology. S.L. worked on software implementation and the implementation of baselines. P.K. contributed to the conceptualization and writing (review and editing). R.Z. and J.S.L. both were responsible for the methodology and writing (review). T.A. handled the implementation of baselines. I.Y. handled data acquisition. B.W. contributed to the methodology, writing (review and editing), and data acquisition. B.M. provided supervision, conceptualization, methodology, and writing (review and editing).
B.W. and B.M. contributed equally as senior authors.
9 Competing Interests
The authors declare no competing interests.